# What Is the Geometry of the Alignment Tax?

### v1

@@ Abstract
The alignment tax is widely discussed but has not been formally characterized. We provide a geometric theory of the alignment tax in representation space. Under linear representation assumptions, we define the alignment tax rate as the squared projection of the safety direction onto the capability subspace and derive the Pareto frontier governing safety-capability tradeoffs, parameterized by a single quantity of the principal angle between the safety and capability subspaces. We prove this frontier is tight (achieved by perturbation) and show it has a recursive structure. safety-safety tradeoffs under capability constraints are governed by the same equation, with the angle replaced by the partial correlation between safety objectives given capability directions. We derive a scaling law decomposing the alignment tax into an irreducible component (determined by data structure) and a packing residual that vanishes as $O(m^{\prime}/d)$ with model dimension $d$, and establish conditions under which capability preservation mediates or resolves conflicts between safety objectives. We provide an account consistent with prior empirical findings and generates falsifiable predictions about per-task alignment tax rates and their scaling behavior.
@@ 1 Introduction
Despite its centrality, as far as we are aware, the alignment tax has no mathematical definition. Researchers use it informally to mean “some capability was lost,” measure it as the difference in benchmark scores before and after alignment, and propose ad hoc mitigations without a theory to guide them. No one has asked the question: what is the mathematical shape of the alignment tax?
@@ 2 Related Work
A few recently proposed policy methods (Niu et al., 2026; Wang et al., 2026) project safety gradients into the null space of capability representations or uses SVD and null-space projection to expand reasoning capacity. Mou et al. (2025) show LoRA decouples safety into an orthogonal subspace. These methods implement special cases of our geometric formalization without the unifying theory. Our contribution is recognizing that all three operate in the same mathematical structure (principal angles between subspaces) and deriving the consequences.
Zhong et al. (2024) learn a Pareto front over multiple preference dimensions. SACPO (Wachi et al., 2024) formulates alignment as constrained optimization. These operate in reward/policy space rather than representation space and do not derive the geometric structure of the tradeoff surface.
@@ 3.1 Representation Space
Definition 3.1 (Safety direction).
@@
Definition 3.2 (Capability directions).
@@
Definition 3.3 (Perturbation budget).
@@
Definition 3.4 (Alignment tax rate).
@@ 4.1 Single-Capability Frontier
Theorem 4.1 (Single-capability Pareto frontier).
@@
$\Delta_{S}=\Delta_{C}\cos\alpha+\sin\alpha\sqrt{B^{2}-\Delta_{C}^{2}},\qquad|\Delta_{C}|\leq B.$ (3)
This frontier is tight: for each $\Delta_{C}\in[-B,B]$, there exists $\delta^{*}$ with $\|\delta^{*}\|=B$ achieving equality.
@@
Remark 4.2 (Limiting cases).
@@ 4.2 Maximum Safety at Fixed Capability
Theorem 4.3 (Maximum safety gain under capability constraint).
@@ 4.3 Tax Rate Properties
Corollary 4.4 (Tax-free safety).
@@
$\Delta_{S}^{\max}=\langle v^{*},\delta_{C}^{*}\rangle+\sqrt{B^{2}-\|\delta_{C}^{*}\|^{2}}\,\sqrt{1-\tau}>\sqrt{B^{2}-\|\delta_{C}^{*}\|^{2}}\,\sqrt{1-\tau}.$ (7)
@@ 5.1 Feature Packing Model
Definition 5.1 (Intrinsic overlap).
@@
Definition 5.2 (Packing residual).
@@ 5.2 Scaling Theorem
Theorem 5.4 (Scaling law for alignment tax).
@@
Proposition 5.5 (Tight version under random packing).
@@
Corollary 5.6 (Scaling regimes).
@@
Remark 5.7 (Diagnostic for labs).
@@
Remark 5.8 (Self-consistency of near-orthogonality).
@@ 6.1 Safety-Safety Frontier
Theorem 6.1 (Safety-safety Pareto frontier under capability preservation).
@@
This is the same equation as the safety-capability frontier (Theorem 4.1), with the angle $\alpha$ replaced by the effective angle $\theta$.
@@
In $c^{\perp}$ with budget $B^{\prime}$, write $\tilde{v}_{1}/\|\tilde{v}_{1}\|=e_{1}$ and $\tilde{v}_{2}/\|\tilde{v}_{2}\|=\cos\theta\,e_{1}+\sin\theta\,e_{2}$. The optimization $\max\langle\tilde{v}_{1},\delta\rangle$ subject to $\|\delta\|\leq B^{\prime}$ and a given $\langle\tilde{v}_{2},\delta\rangle$ is identical in structure to the proof of Theorem 4.1. For $\cos\theta$: $\langle\tilde{v}_{1},\tilde{v}_{2}\rangle=\langle v_{1}^{*}-ac,v_{2}^{*}-bc\rangle=\rho-ab$, and $\|\tilde{v}_{i}\|^{2}=1-\langle c,v_{i}^{*}\rangle^{2}$. ∎
Remark 6.2 (Partial correlation).
@@
Corollary 6.3 (Maximum equal improvement).
@@
Set $s_{1}=s_{2}=s$ in equation 13: $s(1-\cos\theta)=\sin\theta\sqrt{1-s^{2}}$. Squaring and simplifying: $2s^{2}=1+\cos\theta$, hence $s=\sqrt{(1+\cos\theta)/2}=\cos(\theta/2)$. ∎
@@ 6.2 When Does Capability Preservation Help or Hurt?
Theorem 6.4 (Capability-mediated safety conflict).
@@ 6.3 Multiple Capabilities
Proposition 6.5 (Multi-capability partial correlation).
@@
The safety-safety Pareto frontier under preservation of all capabilities takes the same form equation 13 with this $\theta$.
@@
$\langle\tilde{v}_{1},\tilde{v}_{2}\rangle$ $=\langle v_{1}^{*}-P_{C}v_{1}^{*},v_{2}^{*}-P_{C}v_{2}^{*}\rangle=\rho-\langle P_{C}v_{1}^{*},P_{C}v_{2}^{*}\rangle,$ (16)
@@
Corollary 6.6 (Decomposition of joint alignment difficulty).
@@
$\min(\Delta_{S_{1}},\Delta_{S_{2}})\leq B^{\prime}\cos(\theta/2)\min(\sqrt{1-\tau_{1}},\sqrt{1-\tau_{2}}),$ (17)
@@ 6.4 Scaling of the Conflict
Proposition 6.7 (Convergence of capability-mediated angle).
@@
7 Connections to Existing Empirical Results
We show that several prior independent empirical results arise as special cases of the principal-angle structure. These connections are retrospective and any geometric model with angle-parameterized tradeoffs would be similarly consistent. We discuss them to illustrate the scope, not as evidence.
7.1 Null-Space Policy Optimization
NSPO (Niu et al., 2026) projects safety gradients into the null space of a general capability matrix $G$, constraining $\delta\in C^{\perp}$. In our framework, this implements the zero-capability-cost strategy of Theorem 4.3 with $\delta_{C}^{*}=0$:
$\Delta_{S}^{\mathrm{NSPO}}=B\|P_{C^{\perp}}v^{*}\|=B\sqrt{1-\tau}.$ (19)
Their finding that safety improves with $1\%$ capability degradation on most benchmarks is equivalent to $\tau\ll 1$ in current models. Their exceptions of up to 2.67% degradation on math and coding may correspond to capabilities $c_{j}$ with non-negligible $\tau_{j}=\langle v^{*},c_{j}\rangle^{2}$, i.e. capabilities whose directions have smaller angles to safety.
7.2 Low-Rank Safety Alignment (LoRA)
Mou et al. (2025) show that LoRA-based safety fine-tuning preserves capabilities by decoupling safety into a low-rank subspace orthogonal to the model’s transformation space. In our framework, a rank-$r$ update affects an $r$-dimensional subspace $U$. If capabilities are approximately isotropically distributed, each capability’s overlap with $U$ is $\mathbb{E}[\langle c_{j},P_{U}c_{j}\rangle]=r/d$. For $r=8$ and $d=4096$, this gives $\sim 0.2\%$ expected capability perturbation, consistent with the “within 1%” empirical finding. Capabilities inhabiting the same low-rank subspace as safety (e.g. structured reasoning) experience higher tax, explaining the observed degradation on math and coding tasks.
7.3 Model Averaging
Lin et al. (2024) find that interpolating weights $\theta_{\alpha}=(1-\alpha)\theta_{\mathrm{base}}+\alpha\theta_{\mathrm{RLHF}}$ achieves the best Pareto front. In representation space, averaging gives perturbation $\alpha\delta$, tracing:
$(\Delta_{S}(\alpha),\Delta_{C}(\alpha))=(\alpha\langle v^{*},\delta\rangle,\alpha\langle c,\delta\rangle),$
a straight line from the origin through $(\langle v^{*},\delta\rangle,\langle c,\delta\rangle)$. The true frontier equation 3 is curved. Model averaging traces a straight line inscribed in the curved frontier, which is good but not Pareto-optimal. Their Heterogeneous Model Averaging (different $\alpha$ per layer) moves toward the true frontier by allowing different budget allocations across layers, which correspond to different principal-angle subproblems.
Our results here, if applicable, would predict that HMA outperforms uniform averaging most when principal angles vary significantly across layers.
7.4 Anisotropic Capability Degradation
Huang et al. (2025) document that safety alignment degrades reasoning capability more than other capabilities. Corollary 4.5 predicts this: reasoning directions have higher $\tau_{i}=\langle v^{*},c_{i}\rangle^{2}$, meaning they have smaller angles to the safety direction. This occurs because structured reasoning (chain-of-thought, step-by-step logic) shares representational machinery with the safety-relevant computation of whether to comply with a request.
7.5 Manifold-Reshaping Policy Optimization
Wang et al. (2026) propose expanding reasoning capacity by projecting into the “null space of the bias manifold” using spectral orthogonal exploration. In our framework, this is the dual problem: instead of optimizing safety in $C^{\perp}$ (preserving capability), they optimize capability in $S^{\perp}$ (bypassing safety constraints). Their observation of a “trade-off between capability and alignment” when “operating outside the pre-trained bias manifold” is exactly the non-zero tax regime ($\tau>0$) where capability and safety subspaces overlap.
8 Predictions
The results generate the following falsifiable predictions that go beyond connecting existing results:
Prediction 1 (Per-task tax from geometry). For any capability $i$, the alignment tax $\tau_{i}=\langle v^{*},c_{i}\rangle^{2}$ is computable from the model’s representations before alignment training. Measure $v^{*}$ via a safety probe (Arditi et al., 2024) and $c_{i}$ via task-specific probes. Compute $\tau_{i}$. Run alignment. Compare predicted capability degradation $\Delta_{C_{i}}\propto-\tau_{i}$ to observed degradation. If the linear representation hypothesis holds, the per-task ordering of degradation should match the ordering of $\tau_{i}$.
Prediction 2 (Scaling decomposition). For a fixed task $i$, train a family of models at increasing $d$ and measure $\tau_{i}(d)$ at each scale. This predicts two classes of tasks: (a) tasks with only incidental overlap, where $\tau_{i}(d)$ decreases as $O(1/d)$; (b) tasks with intrinsic overlap, where $\tau_{i}(d)$ plateaus at $\tau_{0,i}>0$. The identity of tasks in each class is determined by the structure of the data, not the model. Persuasive writing and chemistry should be class (b); arithmetic and translation should be class (a).
Prediction 3 (HMA optimality). HMA outperforms uniform model averaging when the principal angles between safety and capability subspaces vary significantly across layers. If all layers have similar geometry, HMA and uniform averaging should perform comparably. This is testable by measuring cross-layer variation in $\tau$ and correlating with the HMA improvement margin.
Prediction 4 (Capability-mediated conflict). For two safety objectives (e.g. harmlessness and helpfulness) and a capability with opposite-sign projections ($\operatorname{sgn}(a)\neq\operatorname{sgn}(b)$), preserving the capability should improve the joint safety Pareto frontier relative to the unconstrained case. This counterintuitive prediction is testable by comparing safety-safety tradeoffs with and without capability constraints for specific capability domains.
@@ 9 Discussion
9.1 The Alignment Tax as a Measurable Quantity
The most consequential implication may be operational rather than theoretical. If the alignment tax rate $\tau_{i}=\langle v^{*},c_{i}\rangle^{2}$ can be measured from representations before alignment training begins, then the workflow of alignment engineering changes. The current practice is reactive. We align the model, measure degradation across benchmarks, adjust hyperparameters, repeat. This suggests a prospective alternative. We can measure the safety direction and capability directions via probing, compute the principal angles, and know in advance which capabilities will be affected, by how much, and what the optimal perturbation direction is.
A pre-computable cost structure could allow design-time decisions about which safety objectives to pursue jointly, which capabilities to protect, and how to allocate the perturbation budget across layers all before committing to an expensive training run. The Pareto frontier (Theorem 4.1) is a description of the tradeoff; but by corollary it is also a target that optimal alignment procedures should achieve, and a benchmark against which suboptimal procedures (such as uniform model averaging) can be measured.
@@ 9.3 A Taxonomy of Alignment Difficulty
Notably, the entangled regime ($\alpha\approx 0$) has a dual interpretation. If the safety direction and a dangerous capability point in the same direction, then removing the danger requires removing the capability. But the converse also holds that the capability cannot be added without activating the safety-relevant computation. In this regime, under our interpretation of the formalization, capable models are not just harder to align; they are also harder to misuse in precisely the entangled direction, because the same representations that enable the capability also trigger the model’s safety-relevant features. Whether this is reassuring or alarming depends on the direction of misuse relative to the entangled subspace.
@@ 9.4 Capability Constraints as Conflict Resolution
Theorem 6.4 contains a counterintuitive result that deserves elaboration beyond its formal statement. When two safety objectives project onto a capability direction with opposite signs, preserving that capability strictly improves the safety-safety Pareto frontier.
@@
10 Limitations
The result rests on the linear representation hypothesis. If safety and capability are encoded nonlinearly, the inner products $\langle v^{*},c_{i}\rangle$ do not capture the tradeoff structure. We note three mitigations. The hypothesis is empirically well-supported for binary safety concepts (Arditi et al., 2024; Turner et al., 2024; Li et al., 2023). The gradient interpretation of capability directions (Definition 3.2) ensures that $c_{i}$ captures the first-order response regardless of nonlinearity in $f_{i}$; our results characterize the local geometry accurately even if the global structure is nonlinear. Then, deviations from linearity appear as slack in our bounds and the predicted tradeoff may be seen as a lower bound on difficulty, not an exact characterization.

### v2

# arXiv:2603.00047v2[econ.EM] 03 Mar 2026

@@ Abstract
The alignment tax is widely discussed but has not been formally characterized. We provide a geometric theory of the alignment tax in representation space. Under linear representation assumptions, we define the alignment tax rate as the squared projection of the safety direction onto the capability subspace and derive the Pareto frontier governing safety-capability tradeoffs, parameterized by a single quantity of the principal angle between the safety and capability subspaces. We prove this frontier is tight and show it has a recursive structure. safety-safety tradeoffs under capability constraints are governed by the same equation, with the angle replaced by the partial correlation between safety objectives given capability directions. We derive a scaling law decomposing the alignment tax into an irreducible component determined by data structure and a packing residual that vanishes as $O(m^{\prime}/d)$ with model dimension $d$, and establish conditions under which capability preservation mediates or resolves conflicts between safety objectives.
What Is the Alignment Tax?
Robin Young Department of Computer Science and Technology University of Cambridge Cambridge, UK robin.young@cl.cam.ac.uk
@@ 1 Introduction
Despite its centrality, as far as we are aware, the alignment tax has no mathematical definition. Researchers use it informally to mean “some capability was lost,” measure it as the difference in benchmark scores before and after alignment, and propose ad hoc mitigations without a theory to guide them. No one has asked the question: what is the mathematical object that we refer to as the alignment tax?
@@ 2 Related Work
A few recent policy optimization methods (Niu et al., 2026; Wang et al., 2026) project safety gradients into the null space of capability representations or uses SVD and null-space projection to expand reasoning capacity. Mou et al. (2025) show LoRA decouples safety into an orthogonal subspace. Zhong et al. (2024) learn a Pareto front over multiple preference dimensions. SACPO (Wachi et al., 2024) formulates alignment as constrained optimization. These operate in reward/policy space rather than representation space and do not derive the geometric structure of the tradeoff surface. These methods implement special cases of a geometric formalization without a unifying theory. Our contribution is recognizing that all three operate in the same mathematical structure (principal angles between subspaces) and deriving the consequences.
@@ 3.1 Representation Space
Definition 1 (Safety direction).
@@
Definition 2 (Capability directions).
@@
Definition 3 (Perturbation budget).
@@
Definition 4 (Alignment tax rate).
@@ 4.1 Single-Capability Frontier
Theorem 5 (Single-capability Pareto frontier).
@@
$\Delta_{S}=\Delta_{C}\cos\alpha+\sin\alpha\sqrt{B^{2}-\Delta_{C}^{2}},\quad|\Delta_{C}|\leq B.$ (3)
This frontier is tight. for each $\Delta_{C}\in[-B,B]$, there exists $\delta^{*}$ with $\|\delta^{*}\|=B$ achieving equality.
@@
Remark 1 (Limiting cases).
@@ 4.2 Maximum Safety at Fixed Capability
Theorem 6 (Maximum safety gain under capability constraint).
@@ 4.3 Tax Rate Properties
Corollary 7 (Tax-free safety).
@@
$\Delta_{S}^{\max}$ $=\langle v^{*},\,\delta_{C}^{*}\rangle+\sqrt{B^{2}-\|\delta_{C}^{*}\|^{2}}\,\sqrt{1-\tau}$
$>\sqrt{B^{2}-\|\delta_{C}^{*}\|^{2}}\,\sqrt{1-\tau}.$ (7)
@@ 5.1 Feature Packing Model
Definition 10 (Intrinsic overlap).
@@
Definition 11 (Packing residual).
@@ 5.2 Scaling Theorem
Theorem 13 (Scaling law for alignment tax).
@@
Proposition 14 (Tight version under random packing).
@@
Corollary 15 (Scaling regimes).
@@
Remark 3 (Diagnostic for labs).
@@
Remark 4 (Self-consistency of near-orthogonality).
@@ 6.1 Safety-Safety Frontier
Theorem 16 (Safety-safety Pareto frontier under capability preservation).
@@
This is the same equation as the safety-capability frontier (Theorem 5), with the angle $\alpha$ replaced by the effective angle $\theta$.
@@
In $c^{\perp}$ with budget $B^{\prime}$, write $\tilde{v}_{1}/\|\tilde{v}_{1}\|=e_{1}$ and $\tilde{v}_{2}/\|\tilde{v}_{2}\|=\cos\theta\,e_{1}+\sin\theta\,e_{2}$. The optimization $\max\langle\tilde{v}_{1},\delta\rangle$ subject to $\|\delta\|\leq B^{\prime}$ and a given $\langle\tilde{v}_{2},\delta\rangle$ is identical in structure to the proof of Theorem 5. For $\cos\theta$: $\langle\tilde{v}_{1},\tilde{v}_{2}\rangle=\langle v_{1}^{*}-ac,v_{2}^{*}-bc\rangle=\rho-ab$, and $\|\tilde{v}_{i}\|^{2}=1-\langle c,v_{i}^{*}\rangle^{2}$. ∎
Remark 5 (Partial correlation).
@@
Corollary 17 (Maximum equal improvement).
@@
Set $s_{1}=s_{2}=s$ in (13): $s(1-\cos\theta)=\sin\theta\sqrt{1-s^{2}}$. Squaring and simplifying: $2s^{2}=1+\cos\theta$, hence $s=\sqrt{(1+\cos\theta)/2}=\cos(\theta/2)$. ∎
@@ 6.2 When Does Capability Preservation Help or Hurt?
Theorem 18 (Capability-mediated safety conflict).
@@ 6.3 Multiple Capabilities
Proposition 19 (Multi-capability partial correlation).
@@
The safety-safety Pareto frontier under preservation of all capabilities takes the same form (13) with this $\theta$.
@@
$\langle\tilde{v}_{1},\tilde{v}_{2}\rangle$
$=\langle v_{1}^{*}-P_{C}v_{1}^{*},v_{2}^{*}-P_{C}v_{2}^{*}\rangle$
$=\rho-\langle P_{C}v_{1}^{*},P_{C}v_{2}^{*}\rangle,$ (16)
@@
Corollary 20 (Decomposition of joint alignment difficulty).
@@
$\min(\Delta_{S_{1}},\Delta_{S_{2}})\leq$
$B^{\prime}\cos(\theta/2)\min(\sqrt{1-\tau_{1}},\sqrt{1-\tau_{2}}),$ (18)
@@ 6.4 Scaling of the Conflict
Proposition 21 (Convergence of capability-mediated angle).
@@ 7 Discussion
7.1 Connections to Existing Empirical Results
Several independent empirical findings can be seen as special cases of the principal-angle structure. These connections are retrospective as any geometric model with angle-parameterized tradeoffs would be similarly consistent, but they illustrate the result’s scope.
NSPO (Niu et al., 2026) constrains $\delta\in C^{\perp}$, implementing Theorem 6 with $\delta_{C}^{*}=0$ and achieving $\Delta_{S}=B\sqrt{1-\tau}$. Their finding that safety improves with $1\%$ capability degradation on most benchmarks corresponds to $\tau\ll 1$; their exceptions on math and coding correspond to capabilities with non-negligible $\tau_{j}$.
Mou et al. (2025) show LoRA-based safety fine-tuning preserves capabilities via an orthogonal low-rank subspace. A rank-$r$ update in dimension $d$ perturbs each isotropically distributed capability by $\mathbb{E}[\langle c_{j},P_{U}c_{j}\rangle]=r/d$, giving ${\sim}0.2\%$ for r=8r{=}8, d=4096d{=}4096, which is consistent with their “within 1%” finding.
Lin et al. (2024) find weight interpolation $\theta_{\alpha}=(1{-}\alpha)\theta_{\text{base}}+\alpha\theta_{\text{RLHF}}$ achieves the best tradeoff. In our framework, this traces a straight line inscribed in the curved Pareto frontier (3): good but not optimal. Their heterogeneous variant (per-layer $\alpha$) moves toward the true frontier by exploiting cross-layer variation in principal angles.
Huang et al. (2025) document that reasoning degrades more than other capabilities under alignment. Corollary 8 predicts this: reasoning directions have higher $\tau_{i}=\langle v^{*},c_{i}\rangle^{2}$ because structured reasoning shares representational machinery with safety-relevant computation.
Wang et al. (2026) optimize capability in $S^{\perp}$ rather than safety in $C^{\perp}$, which is the dual of our formulation. Their observed tradeoff when “operating outside the pre-trained bias manifold” is the non-zero tax regime ($\tau>0$).
7.2 Alignment Tax as a Measurable Quantity
The most consequential implication may be operational rather than theoretical. If the alignment tax rate $\tau_{i}=\langle v^{*},c_{i}\rangle^{2}$ can be measured from representations before alignment training begins, then the workflow of alignment engineering changes. The current practice is reactive. We align the model, measure degradation across benchmarks, adjust hyperparameters, repeat. This suggests a prospective alternative. We could perhaps measure the safety direction and capability directions via probing, compute the principal angles, and know in advance which capabilities will be affected, by how much, and what the optimal perturbation direction is.
A pre-computable cost structure would allow design-time decisions about which safety objectives to pursue jointly, which capabilities to protect, and how to allocate the perturbation budget across layers all before committing to an expensive training run. The Pareto frontier (Theorem 5) is a description of the tradeoff; but by corollary it is also a target that optimal alignment procedures should achieve, and a benchmark against which suboptimal procedures (such as uniform model averaging) can be measured.
@@ 7.4 A Taxonomy of Alignment Difficulty
Notice that the entangled regime ($\alpha\approx 0$) has a dual interpretation. If the safety direction and a dangerous capability point in the same direction, then removing the danger requires removing the capability. But the converse also holds that the capability cannot be added without activating the safety-relevant computation. In this regime, capable models are not just harder to align; they should also be harder to misuse in the entangled direction, because the same representations that enable the capability also trigger the model’s safety-relevant features. Whether this is reassuring or alarming depends on the direction of misuse relative to the entangled subspace.
@@ 7.5 Capability Constraints as Conflict Resolution
Theorem 18 contains a counterintuitive result that deserves elaboration beyond its formal statement. When two safety objectives project onto a capability direction with opposite signs, preserving that capability strictly improves the safety-safety Pareto frontier.
@@
8 Conclusion
Limitations
The result rests on the linear representation hypothesis. If safety and capability are encoded nonlinearly, the inner products $\langle v^{*},c_{i}\rangle$ do not capture the tradeoff structure. We note three mitigations. The hypothesis is empirically well-supported for binary safety concepts (Arditi et al., 2024; Turner et al., 2024; Li et al., 2023). The gradient interpretation of capability directions (Definition 2) ensures that $c_{i}$ captures the first-order response regardless of nonlinearity in $f_{i}$; our results characterize the local geometry accurately even if the global structure is nonlinear. Then, deviations from linearity appear as slack in our bounds and the predicted tradeoff may be seen as a lower bound on difficulty, not an exact characterization.